# Space Group Equivariant Crystal Diffusion

### v1

@@ Results.
On MP20, we found SGEquiDiff was competitive on all metrics and notably achieved a 60.5% higher S.U.N. rate than the next best model (Table 1). Unsurprisingly, we found that SGEquiDiff and other space group-constrained models (DiffCSP++, SymmCD) strongly outperformed baselines on the JSD metric for space groups and Wyckoff dimensions. We also show that by increasing the inference-time temperature of SGEquiDiff’s learned categorical distributions over space group, Wyckoff position, element, and the stop token, the U.N. rate and composition diversity metric can be significantly improved at the expense of slightly lower validity rates and higher distributional distances to the test set. Inspecting normalized root mean square Cartesian displacements between as-generated and DFT-relaxed structures, we found all benchmarked models performed similarly, with MatterGen being the best on average (Table 3). We note that MatterGen (44.6M parameters) is significantly larger than SGEquiDiff (5.5M parameters).
We found that SGEquiDiff+LDiff improved some proxy metrics over SGEquiDiff including Wasserstein distances for atomic mass density and number of elements (possibly because the Wyckoff-Element Transformer is not conditioned on noisy lattices in the LDiff variant) as well as the diversity metrics. However, SGEquiDiff yielded a higher validity rate, U.N. rate, and S.U.N. rate, which are more relevant metrics for discovering new materials. Despite stable training, we also found the LDiff variant occasionally generated NaN lattices, which is impossible for SGEquiDiff’s autoregressive lattice sampler due to our imposed constraints. We show examples of S.U.N. crystals from both SGEquiDiff variants in Figs. 4 and 5.
@@
We also benchmarked SGEquiDiff on the more difficult and memory-intensive MPTS52 dataset (Table 2), which most other models do not present results on. Using the same hyperparameters as used on MP20, we found that SGEquiDiff is able to scale to more atoms while maintaining a high rate of valid, unique, and novel generated crystals.
Table 2: Results on the MPTS52 dataset.
Sampling time $\downarrow$ Validity (%) $\uparrow$ U.N. rate (%) $\uparrow$ Distribution distance $\downarrow$ CMD $\downarrow$ Diversity $\uparrow$
(sec / batch) Structure Composition $\mathcal{D}^{\mathrm{MP20}}_{\mathrm{train}}$ $W_{\rho}$ $W_{N_{el}}$ $\mathrm{JSD}_{G}$ $\mathrm{JSD}_{d_{\mathrm{Wyckoff}}}$ Structure Structure Composition
SymmCD (reported) [53] - 90.1 79.2 - - - - - - - -
SGEquiDiff 713 97.24 78.42 90.5 0.9320 0.1725 0.2954 0.1427 0.4323 0.8559 16.17
@@ 6 Conclusion
Due to computational constraints, we were only able to conduct 100 DFT relaxations per model, adding unknown variance to our reported S.U.N. rates. SGEquiDiff’s training data was pre-processed to assign space groups to crystals with nonzero tolerances over atom positions and lattice angles; it is possible that some materials were incorrectly symmetrized to higher symmetry space groups than representative of reality. SGEquiDiff also does not provide guidance for how to synthesize generated materials. Other symmetry groups relevant to crystals were ignored, including magnetic space groups [91], spin space groups [95], and layer groups [23]. Practically relevant deviations of real materials from perfect crystals were not considered, including defects, compositional disorder, surface effects, and interfacial effects. Future work might include application to a broader set of tasks such as those derived from non-scalar properties [99, 57, 92] or materials spectroscopy [54]. Exploring inference-time space group symmetrization may also help unify literature on crystal generative models that are not space group-constrained with those that are.
@@ Appendix A Appendix
Table 3: Normalized root mean square displacements (RMSD) in Angstroms and their standard deviations calculated with pymatgen StructureMatcher between as-generated and DFT-relaxed crystals. RMSD is normalized by $(V/N)^{1/3}$ for unit cell volume $V$ and number of atoms per cell $N$.
CDVAE [97] DiffCSP++ [39] MatterGen [104] SGEquiDiff SGEquiDiff+LatDiff
0.0651$\pm$0.0507 0.0742$\pm$0.0766 0.0368$\pm$0.0457 0.0683$\pm$0.0840 0.0860$\pm$0.0810
@@ A.1 Proofs
Proof. Let $\{R,v\}=g_{x}\in G_{x}$ be a stabilizer group element where $R\in O(3)$ is a point group operation, $v\in[0,1)^{3}$ is a fractional lattice translation, and the action of $g_{x}$ is $R(\cdot)+v$. The definition of a stabilizer group requires that
$Rx+v=x\implies v=0.$
@@
$f(Rx)=f(x).$ (10)
The definition of space group equivariance requires that
@@
$\phi(Rx_{t}+v|x_{0})$ =$=\sum_{\{R_{j},t_{j}\}}\exp\Big{(}\frac{-||Rx_{t}+v-(R_{j}x_{0}+t_{j})||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}(R_{j}x_{0}+t_{j}-Rx_{t}-v)$
=$=\sum_{\{R_{j},t_{j}\}}\exp\Big{(}\frac{-||R\big{(}x_{t}-R^{T}(R_{j}x_{0}+t_{j}-v)\big{)}||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}\Big{[}R\big{(}R^{T}(R_{j}x_{0}+t_{j}-v)-x_{t}\big{)}\Big{]}$
=$=\sum_{\{R_{j},t_{j}\}}\exp\Big{(}\frac{-||x_{t}-R^{T}(R_{j}x_{0}+t_{j}-v)||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}\Big{[}R\big{(}R^{T}(R_{j}x_{0}+t_{j}-v)-x_{t}\big{)}\Big{]}$
=$=\sum_{\{R_{j},t_{j}\}}\exp\Big{(}\frac{-||x_{t}-g^{-1}(R_{j}x_{0}+t_{j})\big{)}||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}\Big{[}R\big{(}g^{-1}(R_{j}x_{0}+t_{j})-x_{t}\big{)}\Big{]}$
=$=\sum_{\{R_{j^{\prime}},t_{j^{\prime}}\}}\exp\Big{(}\frac{-||x_{t}-(R_{j^{\prime}}x_{0}+t_{j^{\prime}})||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}\Big{[}R(R_{j^{\prime}}x_{0}+t_{j^{\prime}}-x_{t})\Big{]}$
@@ A.2 Implementation details
$\nabla_{x_{t}}\log q(x_{t}|x_{0})$ $=\nabla_{x_{t}}\log\Bigg{[}\sum_{t_{L}\in\mathbb{Z}^{3}\cap[-m,m]^{3}}\sum_{\{R_{j},t_{j}\}}\exp\Big{(}\frac{-||x_{t}-(R_{j}x_{0}+t_{j}+t_{L})||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}\Bigg{]}$
$=\frac{\sum_{t_{L}\in\mathbb{Z}^{3}\cap[-m,m]^{3}}\sum_{\{R_{j},t_{j}\}}\exp\Big{(}\frac{-||x_{t}-(R_{j}x_{0}+t_{j}+t_{L})||_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}(R_{j}x_{0}+t_{j}+t_{L}-x_{t})}{\sigma_{t}^{2}q(x_{t}|x_{0})}.$ (15)
For each Wyckoff position $w$, we pre-computed approximate values of $\lambda_{t}=\mathbb{E}_{x_{t}\sim p(x_{t}|x_{0}),x_{0}\sim\mathcal{P}_{w}(\mathrm{Uniform(\cdot)})}\big{[}||\nabla_{x_{t}}p(x_{t}|x_{0})||\big{]}^{-1}$ with Monte Carlo sampling. We denote $\mathcal{P}_{w}(\mathrm{Uniform(\cdot)})$ as the uniform distribution on the Wyckoff shapes (see Sec. 4.1 and A.3) corresponding to Wyckoff position $w$. First, we sampled $s$ points $\{x_{0}^{i}\}_{i=1}^{s}$ from $\mathcal{P}_{w}(\mathrm{Uniform(\cdot)})$. Then, for each sample $x_{0}^{i}$, we sampled Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{t}^{2})$ and reparameterized as $x_{t}^{i}=x_{0}^{i}+\mathcal{P}_{x_{0}^{i}}(\epsilon)$. The approximation of $\lambda_{t}$ for Wyckoff position $w$ in space group $G$ was then computed as
@@ A.4.2 Graph Neural Network
$\mathrm{PlaneWaveEmbedding(\cdot)}$ $\leftarrow\big{[}\cos(2\pi(\cdot)\nu_{1})||\sin(2\pi(\cdot)\nu_{1})||\cos(2\pi(\cdot)\nu_{2})||\sin(2\pi(\cdot)\nu_{2})||...\big{]}$
$h_{i}^{0}$ $\leftarrow\mathrm{MLP}\big{(}e_{a_{i}}||\mathrm{PlaneWaveEmbedding}(x_{i})\big{|}|t_{emb})$
$e_{ij}$ $\leftarrow\mathrm{MLP\big{(}\mathrm{PlaneWaveEmbedding}(x_{j}-x_{i})||\mathrm{RBF}(d_{ij})||\mathrm{Norm}(\mathbf{l})\big{)}}$
$f_{ij}^{l}$ $\leftarrow\mathrm{MLP}\big{(}e_{ij}||h_{i}^{l}||h_{j}^{l}\big{)}$
@@ A.5 Density Functional Theory Calculations
Total energy calculations were performed with density functional theory (DFT) using the Vienna Ab Initio Simulation Package (VASP) [52] with Projector Augmented Wave (PAW) pseudopotentials [8]. The Perdew-Burke-Ernzerhof (PBE) [71] exchange correlation functional was used for structural relaxations. All input parameters, plane wave cutoffs, convergence criteria, and k-point densities were determined by pymatgen MPRelaxSet [67] for apples-to-apples comparison with Materials Project data. Crystal stabilities were determined through convex hull analyses using all competing phases available in the Materials Project v2024.11.14 [37].
A.6 Examples of S.U.N. crystals

### v2

@@ Results.
On MP20, we found SGEquiDiff was competitive on all metrics and notably achieved the highest S.U.N. rate (Table 1). Unsurprisingly, we found that SGEquiDiff and other space group-constrained models (DiffCSP++, SymmCD) strongly outperformed baselines on the JSD metric for space groups and Wyckoff dimensions. We also show that by increasing the inference-time temperature of SGEquiDiff’s categorical distributions over space group, Wyckoff position, element, and the stop token, the U.N. rates and composition diversity metric can be significantly improved at the expense of slightly lower validity rates and higher distributional distances to the test set. Inspecting normalized root mean square Cartesian displacements between as-generated and DFT-relaxed structures, we found all benchmarked models performed similarly, with MatterGen being the best on average (Table 4). We note that MatterGen (44.6M parameters) is significantly larger than SGEquiDiff (5.5M parameters). Examples of S.U.N. crystals from SGEquiDiff are shown in Fig. 4.
@@
We also benchmarked SGEquiDiff on the more difficult and memory-intensive MPTS52 dataset (Table 2), which most other models do not present results on. Using the same hyperparameters as used on MP20, we found that SGEquiDiff was able to scale to more atoms while maintaining a high rate of valid, unique, and novel crystals. SGEquiDiff notably achieved the highest U.N. template rate, composition validity, and Wasserstein distances for density and number of elements. Interestingly, all space group-constrained models significantly outperformed DiffCSP in the validity metrics, highlighting the advantage of enforcing space group symmetry as an inductive bias.
Ablations.
To isolate effects of various components of SGEquiDiff, we performed ablations on the MP-20 dataset. First, we replaced our autoregressive lattice sampler with the DDPM-based [33] sampling of E(3)-invariant lattice matrix representations proposed in DiffCSP++ and used in SymmCD (Table 6). We used the same cosine noise scheduler as DiffCSP++, jointly diffusing lattices and atom coordinates. We term this variant of SGEquiDiff as +LDiff. We observed that +LDiff improved the Wasserstein distance for number of elements (possibly because the conditioning of the Wyckoff-Element transformer on lattices was removed) and CMD of structural fingerprints. However, our autoregressive lattice sampler yielded better structural validity, composition validity, JSD of Wyckoff dimensionalities, and S.U.N. rate. We also observed that in roughly $1\%$ of samples, the +LDiff model produced lattice matrices with negative determinant, resulting in a spurious inversion operation that does not preserve the symmetry of the 22 chiral space groups. We show examples of S.U.N. crystals from +LDiff in Fig. 5
We also replaced our space group and Wyckoff position features with those introduced in SymmCD [55] (Table 5). We found our featurizations yielded better structure validity, composition validity, Wasserstein distance for atomic density, structural divergence, and validation log-likelihoods of ground truth Wyckoff positions. In contrast, the SymmCD features yielded more diverse samples, evidenced by higher U.N. rates and composition diversity.
Since space group symmetrization via Eq. 3 can be applied to any periodic translation invariant model, we replaced our FAENet-inspired [20] graph neural network (GNN) with the EGNN-inspired [80] model (CSPNet) used by other crystal generative models [40, 41, 62]. Under a fixed number of trainable parameters (2.2M), we found our GNN yielded better metrics for structural validity, composition validity, and CMD of structural fingerprints (Table 5).
Crystal structure prediction (CSP).
Consistent with previous works [40, 41, 62], we evaluated SGEquiDiff on its ability to predict crystals from a test set using elemental composition as input. Performance was assessed by calculating match rate (MR) and root mean squared error (RMSE) of atomic fractional coordinates based on a single generated crystal per composition. RMSE was normalized by the average free length per atom. Matching was done with pymatgen StructureMatcher [70] using stol=0.5, angle_tol=10, and ltol=0.3. To adapt SGEquiDiff to this task, we combined the pretrained +LDiff variant of SGEquiDiff with the metric learning-based approach proposed in DiffCSP++ (CSPML) for selecting templates from training data and performing element substitutions to satisfy the given composition. Without CSP-specific training (unlike baselines), SGEquiDiff+LDiff achieved competitive performance (Table 3). We note SGEquiDiff could alternatively be adapted for CSP by reordering the factorization in Eq. 5 or by supplying composition as conditioning during training. While more challenging, learning instead of strictly enforcing composition may be desirable since an arbitrary composition is not guaranteed to host a stable crystal (accordingly, experimentally observed compositions can be weighted averages of multiple phases).
Table 3: Crystal structure prediction. Leveraging the template selection and element substitution approach developed in DiffCSP++ [41], we evaluated SGEquiDiff+LDiff on the crystal structure prediction task introduced by DiffCSP [40].
MP-20 MR (%) $\uparrow$ RMSE $\downarrow$
CDVAE* [100] 33.90 0.1045
DiffCSP* [40] 51.49 0.0631
FlowMM* [62] 61.39 0.0566
DiffCSP++ (+CSPML) [41] 70.58 0.0272
SGEquiDiff (+LDiff,CSPML) 69.42 0.0416
* Uses ground truth number of atoms per unit cell.
We highlight that the current CSP task is slightly misaligned with real-world application. Many existing models assume the number of atoms per unit cell will be known beforehand, which is not necessarily true in practice. Additionally, a single composition can host multiple stable crystal structures depending on experimental conditions. For example, iron can exist as a body-centered cubic (2 atoms per conventional cell), face-centered cubic (4 atoms per conventional cell), or at high pressures, hexagonal close-packed structure (6 atoms per conventional cell). Furthermore, materials discovery campaigns are often aimed at composition spaces with sparse training data, violating the i.i.d. assumption of the MP20 dataset splits [100].
@@ 6 Conclusion
Due to computational constraints, we were only able to conduct 100 DFT relaxations per model, adding unknown variance to our reported S.U.N. rates. SGEquiDiff’s training data was pre-processed to assign space groups to crystals with nonzero tolerances over atom positions and lattice angles; it is possible that some materials were incorrectly symmetrized to higher symmetry space groups than representative of reality. SGEquiDiff also does not provide guidance for how to synthesize generated materials. Other symmetry groups relevant to crystals were ignored, including magnetic space groups [94], spin space groups [98], and layer groups [23]. Practically relevant deviations of real materials from perfect crystals were not considered, including defects, compositional disorder, surface effects, and interfacial effects. Future work might include application to a broader set of tasks such as those derived from non-scalar properties [102, 59, 95] or materials spectroscopy [56]. Furthermore, training non-equivariant models with space group invariant target distributions and/or applying inference-time space group symmetrization may enable a broader diversity of model architectures to generate high symmetry crystals.
@@
Acknowledgments and Disclosure of Funding
We acknowledge helpful discussions with Cindy Zhang, Barry Bradlyn, Alex M. Ganose, and Eric Toberer. This research used the Delta advanced computing and data resource (award OAC 2005572) and the Illinois Campus Cluster, operated by the Illinois Campus Cluster Program in conjunction with the National Center for Supercomputing Applications. The research was supported by the National Science Foundation under Grant Nos. DGE 21-46756 and 2118201.
@@ Appendix A Appendix
Table 4: Normalized root mean square displacements (RMSD) in Angstroms and their standard deviations calculated with pymatgen StructureMatcher between as-generated and DFT-relaxed crystals. RMSD is normalized by $(V/N)^{1/3}$ for unit cell volume $V$ and number of atoms per cell NN.
CDVAE [100] DiffCSP++ [41] MatterGen [107] SGEquiDiff SGEquiDiff+LDiff
[…]
@@ A.1 Proofs
Proof. Let $\{R_{g_{x}},v_{g_{x}}\}=g_{x}\in G_{x}$ be a stabilizer group element where $R_{g_{x}}\in O(3)$ is a point group operation, $v_{g_{x}}\in[0,1)^{3}$ is a fractional lattice translation, and the action of $g_{x}$ is $R_{g_{x}}(\cdot)+v_{g_{x}}$. The definition of a stabilizer group requires that
$R_{g_{x}}x+v_{g_{x}}=x\implies v_{g_{x}}=0.$
@@
$f(R_{g_{x}}x)=f(x).$ (10)
The definition of space group equivariance requires that for any $R\in G$,
@@
[…]
$=\sum_{\{R_{j},t_{j}\}}\exp\Big(\frac{-||x_{t}-g^{-1}(R_{j}x_{0}+t_{j})\big)||_{F}^{2}}{2\sigma_{t}^{2}}\Big)\Big[R\big(g^{-1}(R_{j}x_{0}+t_{j})-x_{t}\big)\Big]$
[…]
@@ A.2 Implementation details
[…]
For each Wyckoff position $w$, we pre-computed approximate values of $\lambda_{t}=\mathbb{E}_{x_{t}\sim p(x_{t}|x_{0}),x_{0}\sim\mathcal{P}_{w}(\mathrm{Uniform(\cdot)})}\big[||\nabla_{x_{t}}\log p(x_{t}|x_{0})||\big]^{-1}$ with Monte Carlo sampling. We denote $\mathcal{P}_{w}(\mathrm{Uniform(\cdot)})$ as the uniform distribution on the Wyckoff shapes (see Sec. 4.1 and A.3) corresponding to Wyckoff position $w$. First, we sampled $s$ points $\{x_{0}^{i}\}_{i=1}^{s}$ from $\mathcal{P}_{w}(\mathrm{Uniform(\cdot)})$. Then, for each sample $x_{0}^{i}$, we sampled Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{t}^{2})$ and reparameterized as $x_{t}^{i}=x_{0}^{i}+\mathcal{P}_{x_{0}^{i}}(\epsilon)$. The approximation of $\lambda_{t}$ for Wyckoff position $w$ in space group $G$ was then computed as
@@ A.4.2 Graph Neural Network
[…]
@@ A.5 Density Functional Theory Calculations
Total energy calculations were performed with density functional theory (DFT) using the Vienna Ab Initio Simulation Package (VASP) [54] with Projector Augmented Wave (PAW) pseudopotentials [8]. The Perdew-Burke-Ernzerhof (PBE) [73] exchange-correlation functional was used for structural relaxations. All input parameters, plane wave cutoffs, convergence criteria, k-point densities were determined by pymatgen MPRelaxSet [69]. For direct comparison with Materials Project (MP) data, we applied the MP’s energy correction scheme for anions and mixing GGA/GGA+U calculations [37, 38]. Crystal stabilities were determined through convex hull analyses using all competing phases available in MP v2022.8.23 [39].
A.5.1 Space group and Wyckoff position features
To represent a space group, we created one-hot features indicating the 6 lattice centering types, 6 crystal families, 32 crystallographic point groups, whether the space group is chiral, and whether the space group is centrosymmetric. Since this representation ignores fractional translations from glide and screw symmetries, we counted the maximum number of collisions between space group representations (16), arbitrarily indexed colliding space groups, and created additional one-hot features to avoid collisions. This yielded $6+6+32+1+1+16=62$ features. While the last 16 features are not physically motivated, they give the model the capacity to differentiate space groups and learn their relationships. These features are more abstracted than those of SymmCD [55], which used binary features indicating the 26 space group symmetry operations, 15 symmetry axes, and 7 lattice systems, totaling $7+26\times 15=397$ features.
To featurize a Wyckoff position, we used our 62 space group features; one-hot features for the 100 unique site symmetry symbols across space groups, the 4 numbers of degrees of freedom in atom coordinates (0, 1, 2, or 3), and the 17 unique Wyckoff multiplicities; and 48 Fourier features of Wyckoff shapes’ midpoints and vertices (see Sec 4.1). This yielded $62+100+4+17+48=231$ Wyckoff position features. To compute Fourier features $\nu\in\mathbb{R}^{48}$, we collected the midpoint and vertices of each Wyckoff shape belonging to the Wyckoff position. We used these points $\{p_{i}\in\mathbb{R}^{3\times 1}\}_{i=1}^{n_{w}}$ to compute the following:
$f$ $=2\pi\cdot\mathrm{\ linspace(start=1,end=32,steps=8)}\in\mathbb{R}^{1\times 8}$
$\nu_{i}$ $=\mathrm{flatten}(p_{i}f)\in\mathbb{R}^{24}$
$\nu$ $=\frac{1}{n_{w}}\sum_{i}^{n_{w}}[\sin(\nu_{i}),\cos(\nu_{i})]\in\mathbb{R}^{48}$
We confirmed that this representation yielded no collisions between featurizations of different Wyckoff positions. Our representation differs from that of SymmCD, which represents each Wyckoff position by one-hot features of the 15 symmetry axes and 13 symmetry operations (totaling $15\times 13=195$ features) corresponding to the Wyckoff position’s stabilizer group.
A.6 Ablation results
Table 5: SGEquiDiff Ablations on MP-20. We replaced our space group and Wyckoff position features with those of SymmCD [55] (+SymmCD features) and our denoising graph neural network with that of DiffCSP [40], DiffCSP++ [41], and FlowMM [62] (+CSPNet). For fair comparison in the latter experiment, we fixed the number of message passing steps (5), the number of Gaussian edge frequencies (96), and the number of trainable parameters (2.2M) by reducing CSPNet’s hidden dimension from 512 to 224. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
Validation Wyckoff Validity (%) $\uparrow$ U.N. rate (%) $\uparrow$ Distribution distance $\downarrow$ CMD $\downarrow$ Diversity $\uparrow$
log-likelihood $\uparrow$ Structure Composition Template Structure $W_{\rho}$ $W_{N_{el}}$ $\mathrm{JSD}_{d_{\mathrm{Wyckoff}}}$ Structure Structure Composition
SGEquiDiff -0.3206(83) 99.25(28) 86.16(32) 18.6(16) 80.0(16) 0.1938(1049) 0.2092(678) 0.0252(90) 0.1718(58) 0.9190(204) 15.64(7)
+SymmCD features -0.4925(30) 98.55(21) 82.82(28) 30.0(21) 84.9(18) 0.4217(1525) 0.1706(279) 0.0395(138) 0.3307(125) 0.9078(52) 16.10(30)
+CSPNet - 99.00(5) - - 80.6(18) - - - 0.1823(21) 0.9252(155) -
Table 6: LDiff ablation on MP-20. We replaced our autoregressive lattice sampling method with joint diffusion of atom coordinates and E(3)-invariant lattice matrix representations proposed in DiffCSP++ [41] (+LDiff). We report metrics of validity for 10,000 sampled crystals (including the percentage of generated lattices with positive determinant, a requirement to enforce symmetry of chiral space groups); uniqueness and novelty for 1,000 valid crystals; diversity and goodness-of-fit distributional distances for valid, unique, and novel crystals; and stability for 100 valid, unique, and novel crystals. Stability is defined as having a DFT-predicted $E_{\mathrm{hull}}$ < 0.1 eV/atom with respect to the Materials Project v2022.8.23. Best result is bolded, second best is underlined. Standard deviations over three training runs are reported in parentheses as errors of rightmost digits.
Time $\downarrow$ Validity (%) $\uparrow$ U.N. rate (%) $\uparrow$ Distribution distance $\downarrow$ CMD $\downarrow$ Diversity $\uparrow$ S.U.N. (%) $\uparrow$
(s / batch) $|L|>0$ Structure Composition Template Structure $W_{\rho}$ $W_{N_{el}}$ $\mathrm{JSD}_{G}$ $\mathrm{JSD}_{d_{\mathrm{Wyckoff}}}$ Structure Structure Composition MP-2024
SGEquiDiff 226(12) 100(0) 99.25(28) 86.16(32) 18.6(16) 80.0(16) 0.1938(1049) 0.2092(678) 0.1733(53) 0.0252(90) 0.1718(58) 0.9190(204) 15.64(7) 25.80
+LDiff 227(12) 98.95(49) 97.02(97) 84.10(63) 20.3(20) 80.3(10) 0.1435(613) 0.1292(86) 0.1847(122) 0.0579(151) 0.1549(72) 0.9474(131) 15.87(17) 13.65
A.7 Examples of S.U.N. crystals